# A Mind Cannot Be Smeared Across Time

### v1

@@ 8 Attribution and Implications
The machine consciousness context puts special weight on attribution/measurement under uncertainty and implementability constraints. The Temporal Gap formalism makes a concrete measurement target of the gap between ingredient coverage and co-instantiation for a grounded conjunction.
8.1 Instrumentable separation metrics
Fix a grounded conjunction $g^{0}$ and an objective trajectory $\tau$ (possibly generated under controlled conditions). Define the minimal horizon needed for ingredient coverage versus co-instantiation:
$w_{\mathrm{ing}}(g^{0},\tau)$ $:=\inf\{\Delta\in\mathbb{N}:\exists t\ \tau^{\Delta,1}(t)\in\mathrm{T}(\Diamond_{\Delta}g^{0})\},$
$w_{\mathrm{co}}(g^{0},\tau)$ $:=\inf\{\Delta\in\mathbb{N}:\exists t\ \tau^{\Delta,1}(t)\in\Diamond_{\Delta}\mathrm{T}(g^{0})\}.$
I use the convention $\inf\emptyset=\infty$ (so these live in $\mathbb{N}\cup\{\infty\}$). The strict Temporal Gap pattern is $w_{\mathrm{ing}}<w_{\mathrm{co}}$ (or $w_{\mathrm{co}}=\infty$).
These are, in principle, instrumentable when the base-layer ingredients are defined in terms of architectural state variables (e.g. active contributor sets). In the broader context of AI consciousness assessment, this kind of architectural metric aligns with the view that behaviour alone may be insufficient and that concrete indicator properties should be checked against candidate theories [undefl].
8.2 Contrast-class experiment
Construct paired systems that are behaviourally matched at a reporting layer (same task performance and report interface) but differ in base-layer concurrency capacity:
  1. 1.
    Synchronous substrate: implement a bounded-latency global tick and allow high concurrency at the relevant scale (capacity $c\geq n$).
  2. 2.
    Sequential substrate: implement the same macro update rule but enforce small $c$ by serialising contributor updates.
Then select candidate contents $l^{m}$ whose groundings include large conjunctions (Section 6), and measure $w_{\mathrm{ing}}$ and $w_{\mathrm{co}}$. StrongSync predicts that phenomenality (if present) tracks bounded $w_{\mathrm{co}}$ at the relevant grounding layer. WeakSync predicts it may track $w_{\mathrm{ing}}$ alone.
8.3 Normative significance under uncertainty
If StrongSync is even plausible, then behavioural equivalence between a synchronous system and a sequential emulation does not settle whether they are phenomenally equivalent. This creates an ethically salient zone of reasonable disagreement about moral status, particularly when a system could have valenced experience. A precautionary stance toward sentience under uncertainty has been defended in nearby contexts [undefi]. Within my framework, one concrete precautionary recommendation is to treat large Temporal-Gap regimes ($w_{\mathrm{co}}\gg w_{\mathrm{ing}}$ or $w_{\mathrm{co}}=\infty$) as a high-uncertainty region for attribution, motivating conservative deployment and experimentation policies until the StrongSync/WeakSync question is better resolved.
@@ 9 Conclusion
I provided a Stack-Theory-compatible Stack-Time Semantics addendum and used it to sharpen the Temporal Gap into a precise algebraic and architectural question. Compositional grounding (Theorem 1) ensures that any higher-layer phenomenal statement has a base-layer grounded conjunction with identical truth conditions. Induced layer time partitions objective time into maximal constant macro-blocks (a change-point quotient) and can be arbitrarily sparse (Proposition 1), so subjective moments need not correspond to single objective instants. Against that backdrop, the Temporal Gap becomes the non-commutation pattern
@@
with strict inclusion in general for $\Delta\geq 1$ and multi-ingredient statements (Theorem 3, Remark 4). This yields a clean separation between ingredient-wise window occurrence and objective co-instantiation.
StrongSync and WeakSync are then competing postulates about which notion constrains phenomenality. Section 7 notes that conscious level often covaries with temporally coordinated, system-level dynamics (synchrony/connectivity/complexity), which tentatively favours StrongSync over WeakSync.
Finally, I proved that a simple architectural invariant (concurrency capacity) can force the strict Temporal Gap pattern for large conjunctions (Theorem 5). If StrongSync is correct and if a phenomenal content’s base grounding requires a conjunction larger than a substrate’s simultaneous causal contribution can support, then that content cannot be phenomenally realised on that substrate. Under StrongSync, software consciousness on a strictly sequential (capacity-$1$) substrate is impossible for any content whose grounding requires two or more simultaneous contributors. Real machines are not perfectly sequential at the physical level. Thus the practical question becomes which physical concurrency scale is relevant for grounding. Consciousness according to Stack Theory requires a “tapestry of valence”, supported by a solid-brained polycomputing architecture, in which each co-instantiated ingredient simultaneously acts upon the others as they act on it, all linked and synchronised. An extreme concurrency capacity may be needed, and still concurrency is only one (important) aspect of this. The central conclusion remains that consciousness attribution is hardware-sensitive, and surface-level observable behavioural equivalence alone cannot wash out the Temporal Gap.

### v2

@@ 8 Attribution and Implications
The Temporal Gap is the question of whether consciousness is affected by the gap between ingredient coverage and co-instantiation for a grounded conjunction.
Fix a grounded conjunction $g^{0}$ and an objective trajectory $\tau$. Define the minimal horizon needed for ingredient coverage versus co-instantiation:
$w_{\mathrm{ing}}(g^{0},\tau)$ $:=\inf\{\Delta\in\mathbb{N}:\exists t\ \tau_{\Delta}(t)\in\mathrm{T}(\Diamond_{\Delta}g^{0})\},$
$w_{\mathrm{co}}(g^{0},\tau)$ $:=\inf\{\Delta\in\mathbb{N}:\exists t\ \tau_{\Delta}(t)\in\Diamond_{\Delta}\mathrm{T}(g^{0})\}.$
The strict Temporal Gap pattern is $w_{\mathrm{ing}}<w_{\mathrm{co}}$ (or $w_{\mathrm{co}}=\infty$). These are instrumentable when base-layer ingredients are defined in terms of architectural state variables. This kind of architectural metric aligns with the view that behaviour alone may be insufficient for consciousness assessment and that concrete indicator properties should be checked against candidate theories [14].
A concrete test follows: construct paired systems that are behaviourally matched at a reporting layer but differ in base-layer concurrency capacity (one synchronous with $c\geq n$, one sequential with small $c$). Select candidate contents whose groundings include large conjunctions, and measure $w_{\mathrm{ing}}$ and $w_{\mathrm{co}}$. Chord predicts that phenomenality tracks bounded $w_{\mathrm{co}}$; Arpeggio predicts it may track $w_{\mathrm{ing}}$ alone.
If Chord is even plausible, behavioural equivalence between a synchronous system and a sequential emulation does not settle phenomenal equivalence. This creates an ethically salient zone of reasonable disagreement about moral status, particularly when a system could have valenced experience. A precautionary stance toward sentience under uncertainty has been defended in nearby contexts [11]. One concrete recommendation is to treat large Temporal-Gap regimes ($w_{\mathrm{co}}\gg w_{\mathrm{ing}}$ or $w_{\mathrm{co}}=\infty$) as a high-uncertainty region for attribution, motivating conservative deployment policies until the Temporal Gap is better resolved.
@@ 9 Conclusion
I provided a Stack-Theory-compatible Stack-Time Semantics addendum and used it to make the Temporal Gap into a measurable algebraic and architectural question. Compositional grounding (Theorem 1) ensures any higher-layer phenomenal statement has a base-layer grounded conjunction with identical truth conditions. Induced layer time partitions objective time into maximal constant macro-blocks and can be arbitrarily sparse (Proposition 1), so subjective moments need not correspond to single objective instants. The Temporal Gap becomes a question of the non-commutation pattern
@@
with strict inclusion in general for $\Delta\geq 1$ and multi-ingredient statements (Theorem 3). This yields a clean separation between ingredient-wise window occurrence and objective co-instantiation.
Chord and Arpeggio are competing postulates about which notion constrains phenomenality. Section 7 notes that conscious level covaries with temporally coordinated dynamics, tentatively favouring Chord.
A simple architectural invariant (concurrency capacity) can force the strict Temporal Gap pattern for large conjunctions (Theorem 4). Under Chord, software consciousness on a strictly sequential substrate is impossible for any content whose grounding requires two or more simultaneous contributors. Real machines are not perfectly sequential at the physical level, so the practical question becomes which physical concurrency scale is relevant for grounding. Consciousness according to Stack Theory requires a “tapestry of valence” supported by a solid-brained polycomputing architecture in which each co-instantiated ingredient simultaneously acts upon the others, all linked and synchronised (a condition formally included in Chord in [10]). Extreme concurrency capacity may be needed, and still concurrency is only one aspect of this. The central conclusion remains that consciousness attribution is hardware-sensitive, and surface-level behavioural equivalence alone cannot wash out the Temporal Gap.